### v2

@@ 4.2 Results
Multilingual performance. As shown in Table 1, the SEA-HELM benchmark performance demonstrates that our instruct models, Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, attain competitive performance in SEA languages, with Gemma-SEA-LION-v3-9B-IT achieving one of the highest average performances. Both Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT outperform other SEA languages-focused LLMs, such as Sailor2-8B-Chat and SEALLMs-v3-7B-Chat, with an average score of 69.35 across all the languages covered by the SEA-HELM benchmark apart from the SEA-MTBench tasks.
English performance. The Open LLM Leaderboard performance is shown in Table 2. Both Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT performed competitively in English language, math, and reasoning tasks, with Gemma-SEA-LION-v3-9B-IT achieving the highest average score of 35.43.
@@ 4.3 Performance Analysis
Continued Pre-Training The CPT stage is primarily focused on gaining SEA languages capabilities and knowledge. For the purpose of comparison against base and CPT models, we observed a 6.05 and 7.19 average SEA-HELM performance increase over the Meta-Llama-3.1-8B and Gemma-2-9B for Llama-SEA-LION-v3-8B and Gemma-SEA-LION-v3-9B, respectively. We observed a much larger average increase with instruction following capabilities in particular, which we attribute to the fact that our CPT models are trained from the instruct models rather than from the base models. Both CPT models also managed to perform competitively against the Meta-Llama-3.1-8B and Gemma-2-9B base models on the Open LLM Leaderboard benchmarks. This indicates that our choice of retraining with a proportion of 25% English tokens has been beneficial in mitigating catastrophic forgetting which has been shown to stem from CPT (Zheng et al., 2024a).
As shown in Table 1, we chose Gemma since it is the most performant on multilingual benchmarks. However, we also show that our framework generalizes for every LLM by applying our framework on Llama3.1; although the performance of Llama3.1 is lower than Qwen or Sailor, we can still improve it to outperform all of them. Note that we have shown the full performance score of our CPT models and other base models in Appendix A.1.
@@
Final merge: Combining aligned models To compensate for the capability degradation in the previous steps, we merge Merge-2-Llama and Merge-1-Gemma with Aligned-SimPO-Llama and Aligned-SimPO-Gemma and various open sourced pre-trained models describe in sections 3.2.1 and 3.2.2 for their respective model families. For Llama-SEA-LION-v3-8B-IT, we observed a significant increase in average SEA-HELM performance (61.84) from the alignment stage (51.30), mainly from the increase in performance for the core tasks in SEA-HELM. This performance increase demonstrates the value of empirical selection of pre-trained models to be merged in based on each model’s strengths and weaknesses to produce a far superior model. For Gemma-SEA-LION-v3-9B-IT, it easily achieves higher performance compared to the Llama-SEA-LION-v3-8B-IT with fewer post training steps. We attribute this performance to the high performance of the base Gemma 2 model and also to the larger vocabulary size which have been demonstrated (Takase et al., 2024) to produce better models.
@@ 5 Conclusion
Despite the sizable population and language diversity in Southeast Asia, there remains a scarcity of resources and accurate linguistic and cultural representation with open source LLMs. In this paper, we introduce Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, two multilingual LLMs comprehensively trained to achieve state-of-the-art performances in SEA languages, based on the Llama and Gemma family of LLMs. SEA-LION represents the next advancement in the development of LLMs that explicitly supports SEA languages. Both models are fully open-source and available for commercial use to increase accessibility and innovation in multilingual LLMs in Southeast Asia.
@@ Appendix A Appendix
A.1 Benchmarks for Continued Pretrained Models
SEA-HELM
NLU, NLG, NLR, NLI Instruction Following
Models Average ID VI TH TA ID VI TH
Meta-Llama-3.1-8B 35.37 42.33 40.67 35.13 38.88 16.19 19.05 9.00
SeaLLMs-v3-7B 37.04 44.79 48.29 43.53 27.45 26.67 35.24 26.00
Gemma-2-9B 41.48 47.65 43.28 42.00 53.26 4.76 3.81 10.00
Qwen2.5-7B 41.98 51.63 52.17 46.55 36.60 31.43 36.19 30.00
Sailor2-8B 42.62 53.23 47.33 46.64 45.04 30.48 30.48 35.00
Llama-SEA-LION-v3-8B 41.42 44.98 46.25 42.79 43.03 25.71 32.38 23.00
Gemma-SEA-LION-v3-9B 48.67 57.16 49.39 47.16 60.56 25.71 20.00 27.00
Open LLM Leaderboard
Models Average MMLU-PRO BBH GPQA MATH Lvl 5 IFEval (EN) MUSR
Meta-Llama-3.1-8B 13.9 24.95 25.29 6.32 5.14 12.7 8.98
Sailor2-8B 17.71 25.74 27.62 4.87 7.02 21.95 19.03
Gemma-2-9B 21.15 34.48 34.1 10.51 13.14 20.4 14.3
SeaLLMs-v3-7B 24.00 35.71 34.57 9.28 18.81 32.94 12.68
Qwen2.5-7B 24.99 37.39 35.81 9.96 18.88 33.74 14.14
Llama-SEA-LION-v3-8B 16.61 27.6 26.04 7.49 9.89 16.56 12.07
Gemma-SEA-LION-v3-9B 22.41 32.78 37.24 10.29 9.89 30.12 14.11

### v3

@@ 5.2 Instruction Fine-tuning Results
SEA-HELM. As shown in Table 3, the SEA-HELM benchmark performance demonstrates that our instruct models, Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, attain competitive performance in SEA languages, with Gemma-SEA-LION-9B-IT achieving one of the highest average performances. Moreover, we significantly improve the performance of Llama-3.1-8B-Instruct from 41.48 to 61.84 using Llama-SEA-LION-8B-IT, while Gemma-SEA-LION-9B-IT achieves 14.02 improvement points compared to Gemma-2-9B-IT. Both Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT outperform other SEA languages-focused LLMs, such as Sailor2-8B-Chat and SEALLMs-v3-7B-Chat, with an average score of 69.35 across all the languages covered by the SEA-HELM benchmark, apart from the SEA-MTBench tasks. This conforms with the previous results on the CPT models (Section 5.1) that our CPT model performs the best on SEA languages, resulting in the best performer in this experiment.
SEACrowd. Other than evaluating on some SEA languages like SEA-HELM, we also evaluated our model compared to competitors on 31 SEA indigenous languages using SEACrowd-NLU. Note that, for this study, we use only the best settings of our models from the previous experiment (Table 3). As shown in Table 4, we observe a state-of-the-art result from Gemma-SEA-LION-9B-IT by achieving 64.13 points on the NLU benchmark, while Llama-SEA-LION-8B-IT improves its baseline from 49.94 to 55.10 points. Moreover, the results from Figure 3 also emphasize the robustness of our model by reaching more than 80 points on this benchmark, while SeaLLMs and Llama-3.1 have only a few cases where the performance exceeds 80 points. These results emphasize the robustness of our models by achieving the state-of-the-art with a model parameter less than 10B on SEA benchmarks, including both traditional classical NLP benchmark (SEACrowd-NLU) and modern LLM benchmark (SEA-HELM).
English performance. We also evaluate the performance of a widely used language, English, to observe a difference between the results of SEA and English. The Open LLM Leaderboard performance is shown in Table 5. Both Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT performed competitively in English language, math, and reasoning tasks, with Gemma-SEA-LION-9B-IT achieving the highest average score of 35.43. Moreover, we notice that the SEA models (Sailor and SeaLLMs) failed to perform on the English dataset. This might be because these models are optimized for SEA languages during supervised fine-tuning, and English performance decreased as a result. In contrast, our models balance the performance between SEA and English knowledge, resulting in a high score for all benchmarks.
@@ 5.3 Performance Analysis
In this study, we discuss the performance improvement in each design decision of our models (Tables 3 and  5) as follows.
@@
Final merge: Combining aligned models To compensate for the capability degradation in the previous steps, we merge Merge-2-Llama and Merge-1-Gemma with Aligned-SimPO-Llama and Aligned-SimPO-Gemma and various open sourced pre-trained models describe in sections 3.2.1 and 3.2.2 for their respective model families. For Llama-SEA-LION-8B-IT, we observed a significant increase in average SEA-HELM performance (61.84) from the alignment stage (51.30), mainly from the increase in performance for the core tasks in SEA-HELM. This performance increase demonstrates the value of empirical selection of pre-trained models to be merged in based on each model’s strengths and weaknesses to produce a far superior model. For Gemma-SEA-LION-9B-IT, it easily achieves higher performance compared to the Llama-SEA-LION-8B-IT with fewer post training steps. We attribute this performance to the high performance of the base Gemma 2 model and also to the larger vocabulary size which have been demonstrated (Takase et al., 2024) to produce better models.
6 Related Works
Recently, researchers have proposed large language models that support multilingual settings. Llama Dubey et al. (2024) is the prior effort to release an open-source large language model for the research community to develop their own models. Then, Qwen Yang et al. (2024a) and Gemma Rivière et al. (2024) introduced open-source LLMs that perform comparably or better than Llama with a larger amount of training data and many supported languages for these recent models. Massively multilingual open-source models like Bloom Scao et al. (2022) and Aya Üstün et al. (2024) also support a very wide range of languages, including some SEA languages. Although these models demonstrate a robust performance in English benchmarks, they mostly underperformed on SEA benchmarks that tested for SEA languages, SEA knowledge and cultural understanding Lovenia et al. (2024); Susanto et al. (2025), presumably due to a lack of language support for certain SEA languages or cultures.
In the SEA community, many works propose a large language model that is designed specifically for SEA languages by adding more SEA tokens in the training process, such as SeaLLMs Nguyen et al. (2024) and Sailor Sailor2 Team (2024). However, the performance of these models is robust only on in-domain datasets or favors only some tasks (i.e., classical NLP datasets). This is because the design choice in the pre-training or fine-tuning of these models is not well studied, e.g., performing a single SFT step with low-quality datasets written in some SEA languages, resulting in a slight improvement on SEA benchmarks. To create a robust SEA LLM, we need to carefully balance language representation and design both pre-training and post-training (i.e., SFT, alignment, and model merging) for SEA contexts.
@@ 7 Conclusion
Despite the sizable population and language diversity in Southeast Asia, there remains a scarcity of resources and accurate linguistic and cultural representation with open-source LLMs. In this paper, we introduce Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, two multilingual LLMs comprehensively trained to achieve state-of-the-art performances in SEA languages, based on the Llama and Gemma family of LLMs. SEA-LION represents the next advancement in the development of LLMs that explicitly supports SEA languages. Both models are fully open-source and available for commercial use to increase accessibility and innovation in multilingual LLMs in Southeast Asia. We will make our resources publicly available — including the dataset, training scripts, training checkpoints, and all fine-tuned models, even those that achieve state-of-the-art performance on the benchmarks — to establish solid baselines, ensure reproducibility, and support future research focused on culturally and professionally relevant SEA applications.
Limitation
Although we propose the state-of-the-art SEA LLMs, we found that the benchmark might not cover all the properties and languages we want to evaluate. For example, SEA-HELM is a robustness benchmark, but only covers four languages. SEACrowd is a benchmark that covers all SEA languages, but it is only classical NLP datasets (no chat or instruction following datasets). We require a more holistic SEA benchmark that covers LLM-specific tasks written in all SEA languages. However, with the current evaluation design choice, these benchmarks are the best design choice for current SEA research works.
Moreover, we conduct experiments using only 8 and 9 billion parameter models. We argue that this is the most commonly used model size in real-world scenarios. In addition, our method can and should also work with a higher or smaller model size since our proposed technique does not rely on the model size, as we demonstrated by applying the SFT and alignment techniques on both Llama and Gemma models.
@@ Appendix A Appendix
A.1 Continued pre-training (CPT) data
Existing data: We utilize existing datasets as shown in Table 6 (HuggingFace Datasets).
Other data: As shown in Table 6 (the other data section), the listed datasets contain data from a diverse range of domains, including news, books, articles, poems, etc.
Continued Pre-training Data
Source (HuggingFace Datasets) Languages Size (Billions of Tokens)
bigcode/the-stack-v2-dedup CODE 40
allenai/dolma EN 37.5
HuggingFaceFW/fineweb-edu EN 7.5
aisingapore/SEA-PILE-v1 SEA 47.58
aisingapore/SEA-PILE-v2 ID 7
Source (Others) Languages Size (Billions of Tokens)
VinBigData VI 16
WangChanBERTa TH 8.5
Others - EN EN 5
Others - SEA SEA 30.92
Table 6: List of datasets for the continued pre-training stage.
A.2 Stage 1 IFT data
Stage 1 IFT Datasets
Source (HuggingFace Datasets) Languages Size
BAAI/Infinity-Instruct EN 7,449,106
nvidia/OpenMathInstruct-2 EN 2,000,000
Table 7: List of datasets for Stage-1-IFT. For BAAI/Infinity-Instruct dataset, any conversation that originally ended with a user turn has had that last turn removed.
A.3 Stage 2 IFT data
Existing data: We utilize existing datasets as shown in Table 9 (HuggingFace Datasets).
Synthetic data: As shown in Table 9 (the generated part), we describe how to formulate synthetic data as follows
  • qwen_gemma_synthetic datasets are generated first in English with Qwen 32B, utilizing an approach similar to Magpie. Instructions are then translated into the target language with Gemma 2 27B.
  • llama_gemma_synthetic datasets are generated first in English with Llama 3.1 70B, utilizing an approach similar to Magpie Xu et al. (2024). Instructions are then translated into the target language with Gemma 2 27B.
  • gemma_synthetic datasets are generated directly with Gemma 2 27B using Magpie Xu et al. (2024).
  • sea_multilingual_systemchat is a synthetic dataset translated with Gemma 2 27B from the English systemchat dataset.
  • rewritten_oasst is a dataset rewritten with Gemma 2 27B based on the English OASST dataset.
  • rewritten_helpsteer is a dataset rewritten with Gemma 2 27B based on the English Helpsteer dataset.
A.4 Helpfulness and preference alignment data
As shown in Table 8, we use the princeton-nlp/gemma2-ultrafeedback-armorm as the source of the alignment data. We then further re-scored with the reward model, nvidia/Llama-3.1-Nemotron-70B-Reward to create the SEA version. In particular, generated-gemma2-27b-seapref-nemotron-70b takes prompts from seald, wangchan_thaiinstruct, and additional hand-written Southeast Asian cultural prompts collected from native speakers and then generates responses (with a varying temperature) from them with Gemma 2 27B. The responses are then scored with nvidia/Llama-3.1-Nemotron-70B-Reward, with the top-scoring response selected as chosen and vice versa, similar to princeton-nlp/gemma2-ultrafeedback-armorm.
Preference Data
Source (HuggingFace Datasets) Languages Size
princeton-nlp/gemma2-ultrafeedback-armorm EN 61,510
Source (Generated) Languages Size
generated-gemma2-27b-seapref-nemotron-70b SEA 5,511
Table 8: List of preference datasets used for the alignment stage.
Stage 2 IFT Datasets
Source (HuggingFace Datasets) Languages Size
BAAI/Infinity-Instruct^* EN 1,456,927
HuggingFaceTB/smoltalk EN 409,537
allenai/tulu-3-sft-personas-math EN 149,960
parinzee/seed-free-synthetic-instruct-thai-v1 TH 118,898
HuggingFaceTB/smoltalk EN 96,356
HuggingFaceTB/smoltalk EN 83,144
arcee-ai/EvolKit-75K EN 74,174
AI-MO/NuminaMath-TIR EN 72,441
Post-training-Data-Flywheel/AutoIF-instruct-61k EN 61,492
argilla/ifeval-like-data EN 56,339
HuggingFaceTB/smoltalk EN 53,342
ai2-adapt-dev/tulu_v3.9_wildjailbreak_decontaminated_50k EN 50,000
ai2-adapt-dev/tulu_v3.9_synthetic_finalresp_wildguardmixtrain_decontaminated_50k EN 50,000
allenai/tulu-3-sft-personas-math-grade EN 49,980
allenai/tulu-3-sft-personas-code EN 34,999
HuggingFaceTB/smoltalk EN 34,424
allenai/tulu-3-sft-personas-instruction-following EN 29,980
airesearch/WangchanThaiInstruct TH 25,014
allenai/tulu-3-sft-personas-algebra EN 20,000
arcee-ai/EvolKit-20k-vi VI 15,378
allenai/coconot EN 10,983
ai2-adapt-dev/tulu_v3.9_sciriff_10k EN 10,000
Source (Generated) Languages Size
qwen_gemma_synthetic_tamil TA 480,000
qwen_gemma_synthetic_thai TH 480,000
qwen_gemma_synthetic_indonesian ID 465,019
qwen_gemma_synthetic_vietnamese VI 465,019
gemma_synthetic_indonesian ID 458,149
gemma_synthetic_filipino TL 455,093
gemma_synthetic_viet VI 291,576
gemma_synthetic_tamil TA 276,314
gemma_synthetic_thai TH 186,339
gemma_synthetic_javanese JV 110,000
gemma_synthetic_sudanese SU 110,000
llama_gemma_synthetic_thai TH 88,920
llama_gemma_synthetic_tamil TA 88,920
llama_gemma_synthetic_vietnamese VI 88,920
llama_gemma_synthetic_javanese JV 88,920
llama_gemma_synthetic_indonesian ID 88,920
llama_gemma_synthetic_filipino TL 80,000
enrich_27k SEA 27,463
sea_multilingual_systemchat SEA 1,903
rewritten_oasst SEA 841
rewritten_helpsteer SEA 838
Table 9: List of datasets for Stage-2-IFT.